# Deep Improvement Supervision

### v2

@@ 6 Discussion
Potential algorithmic improvements. A key limitation of the current implementation is the use of a fixed number of supervision steps for every ARC task. However, task complexity varies significantly; some tasks may benefit from a higher number of denoising steps, while others require fewer. This observation aligns with findings from the original TRM paper, which highlighted the significant contribution of the halting mechanism to final performance. Therefore, explicitly predicting the necessary number of denoising steps for each task could potentially enhance overall model efficiency and accuracy.
@@
Another promising direction for technical improvement involves adopting a discrete latent space. This approach has been successfully utilized in deep learning architectures such as the Dreamer model (hafner2019dream) and VQ-VAE (razavi2019generating), where latent spaces have proven to be robust and scalable for generative tasks.
@@
Alternative Improvement Generators. As outlined in §4.3, there are several viable methods for generating intermediate steps. While the discrete diffusion prior (Figure 7) serves as the primary source in this work, our framework is designed to support various step-generation approaches.
@@
7 Conclusion
The key contribution is Deep Improvement Supervision, builds on this insight by introducing a structured, stepwise training regime. DIS provides intermediate targets through a discrete diffusion process, transforming the challenging problem of long-term credit assignment into a more tractable supervised learning task (ho2022classifier; frans2025diffusion). This approach not only simplifies training by eliminating the need for learned halting mechanisms but also enhances efficiency, reducing the number of forward passes by 18x with high accuracy.
Appendix A Analysis
Figure 7: The linear corruption process is shown over six steps, from the initial input at time $t=0$ to the target at time $t=6$. A single training sample is illustrated per task.
Notation and Definitions. Let $\mathbf{x}$ be the input context. At any reasoning step $n$, the TRM produces two logit vectors over the vocabulary $\mathcal{V}$:
  • $\ell_{u}^{n}\in\mathbb{R}^{|\mathcal{V}|}$: the reference logits (the state before the current reasoning step),
  • $\ell_{c}^{n}\in\mathbb{R}^{|\mathcal{V}|}$: the conditional logits (the state after the current reasoning step).
We define the Reasoning Residual as
$\Delta\ell^{n}\;=\;\ell_{c}^{n}-\ell_{u}^{n}.$
For a guidance scale $w\geq 0$, define the guided logits $\ell_{w}^{n}:=\ell_{u}^{n}+w\,\Delta\ell^{n}$ and the corresponding policy
$\pi_{w}(a)\;=\;\frac{\exp\!\big(\ell_{u}^{n}[a]+w\,\Delta\ell^{n}[a]\big)}{Z(w)},\qquad Z(w)\;=\;\sum_{k\in\mathcal{V}}\exp\!\big(\ell_{u}^{n}[k]+w\,\Delta\ell^{n}[k]\big).$
Let $y^{\star}$ denote the index of the ground‑truth correct token at the current position. In the proofs below, we fix a step $n$ and drop the superscript $n$ when unambiguous. We first show the condition under which increasing $w$ strictly reduces the loss.
Proposition A.1 (Advantage Margin Condition).
Let $\mathcal{L}(w)=-\log\pi_{w}(y^{\star})$ be the cross‑entropy loss of the correct token $y^{\star}$. Then $\frac{d}{dw}\mathcal{L}(w)<0$ if and only if
$\Delta\ell[y^{\star}]\;>\;\mathbb{E}_{a\sim\pi_{w}}\big[\Delta\ell[a]\big].$
Moreover,
$\frac{d}{dw}\mathcal{L}(w)=\mathbb{E}_{a\sim\pi_{w}}[\Delta\ell[a]]-\Delta\ell[y^{\star}],\qquad\frac{d^{2}}{dw^{2}}\mathcal{L}(w)=\mathrm{Var}_{a\sim\pi_{w}}\!\big[\Delta\ell[a]\big]\geq 0.$
Proof.
Write $\ell_{w}=\ell_{u}+w\,\Delta\ell$.
$\mathcal{L}(w)$ $=-\log\left(\frac{\exp(\ell_{w}[y^{\star}])}{Z(w)}\right)\;=\;-\ell_{w}[y^{\star}]+\log Z(w)$
$=-\big(\ell_{u}[y^{\star}]+w\,\Delta\ell[y^{\star}]\big)+\log\sum_{k\in\mathcal{V}}\exp\big(\ell_{u}[k]+w\,\Delta\ell[k]\big).$
The derivative of the linear term is $\Delta\ell[y^{\star}]$. Let $E_{k}=\exp(\ell_{u}[k]+w\,\Delta\ell[k])$, so $Z(w)=\sum_{k}E_{k}$. Then
$\frac{d}{dw}\log Z(w)=\frac{1}{Z(w)}\sum_{k}E_{k}\,\Delta\ell[k]=\sum_{k}\frac{E_{k}}{Z(w)}\,\Delta\ell[k]=\mathbb{E}_{a\sim\pi_{w}}\!\big[\Delta\ell[a]\big].$
Combining the two derivatives yields
$\frac{d}{dw}\mathcal{L}(w)\;=\;-\Delta\ell[y^{\star}]+\mathbb{E}_{a\sim\pi_{w}}\!\big[\Delta\ell[a]\big],$
which is negative exactly when the stated inequality holds. Differentiating once gives $\frac{d^{2}}{dw^{2}}\mathcal{L}(w)=\mathrm{Var}_{a\sim\pi_{w}}[\Delta\ell[a]]\geq 0$. ∎
Thus, reasoning (the residual $\Delta\ell$) helps iff it boosts the correct class $y^{\star}$ more than it boosts a typical alternative (the policy‑weighted average).
A.1 Global Convergence
While Proposition A.1 guarantees a local improvement as $w$ increases, we also analyze the sequential process with a Lyapunov‑style argument.
Proposition A.2 (Discrete Lyapunov Contraction).
Let $d(\hat{y},y^{\star})$ be a bounded distance metric (e.g., Hamming distance) between a predicted sequence $\hat{y}$ and the ground truth $y^{\star}$. Define the potential at step $n$ as $V_{n}:=d(\mathrm{argmax}(\ell_{c}^{n}),y^{\star})$. Assume:
  1. 1.
    Strictly Improving Targets. A generator $\Phi$ yields $\mathbf{y}^{\dagger}_{0},\dots,\mathbf{y}^{\dagger}_{N}$ with
[…]
  2. 2.
    Training Minimization. The model minimizes $\mathcal{L}_{n}=\mathrm{CE}(\ell_{c}^{n},\mathbf{y}^{\dagger}_{n})$ so that $\mathrm{argmax}(\ell_{c}^{n})=\mathbf{y}^{\dagger}_{n}$ with high probability.
Then $V_{n}-V_{n-1}\leq-\delta<0$. Consequently, $V_{n}$ decreases monotonically and hits 0 in at most $\lceil V_{0}/\delta\rceil$ steps.
Proof.
By Assumption 2, $\mathrm{argmax}(\ell_{c}^{n})\approx\mathbf{y}^{\dagger}_{n}$, hence $V_{n}\approx d(\mathbf{y}^{\dagger}_{n},y^{\star})$ and $V_{n-1}\approx d(\mathbf{y}^{\dagger}_{n-1},y^{\star})$. Assumption 1 gives $d(\mathbf{y}^{\dagger}_{n},y^{\star})\leq d(\mathbf{y}^{\dagger}_{n-1},y^{\star})-\delta$, therefore $V_{n}\leq V_{n-1}-\delta$. Since $V_{n}\geq 0$, the bound on the hitting time follows. ∎
Proposition A.3 (Guaranteed Improvement).
Assume the target generator $\Phi$ provides strictly improving targets such that the likelihood ratio $\log\frac{P(\mathbf{y}^{\dagger}_{n})}{P(\mathbf{y}^{\dagger}_{n-1})}>0$ (with respect to a fixed scoring distribution $P$ over sequences). Then minimizing the sequential loss $\mathcal{L}_{\text{DIS}}$ drives the expected Advantage Margin to be positive:
$\mathbb{E}\!\left[\;\Delta\ell[y^{\star}]\;-\;\mathbb{E}_{a\sim\pi_{w}}\Delta\ell[a]\;\right]\;>\;0,\qquad\text{for any fixed }w\geq 0.$
Proof.
1. Optimal‑logit approximation. Assume sufficient capacity and that $\mathcal{L}_{\text{DIS}}$ is minimized. Then for some constants $C_{1},C_{2}$ (independent of $a$),
$\ell_{c}^{n}\;\approx\;\log P(\cdot\mid\mathbf{y}^{\dagger}_{n})+C_{1},\qquad\ell_{u}^{n}\;\approx\;\log P(\cdot\mid\mathbf{y}^{\dagger}_{n-1})+C_{2}.$
2. Residual as a likelihood ratio. Taking the difference eliminates constants:
$\Delta\ell\;=\;\ell_{c}^{n}-\ell_{u}^{n}\;\approx\;\log\frac{P(\cdot\mid\mathbf{y}^{\dagger}_{n})}{P(\cdot\mid\mathbf{y}^{\dagger}_{n-1})}.$
3. Margin at the ground‑truth class. For the correct token $y^{\star}$,
$\Delta\ell[y^{\star}]\;\approx\;\log\frac{P(y^{\star}\mid\mathbf{y}^{\dagger}_{n})}{P(y^{\star}\mid\mathbf{y}^{\dagger}_{n-1})}\;>\;0$
by the “strictly improving targets” premise.
4. Policy‑weighted average. As the targets sharpen toward $y^{\star}$, probability mass shifts into $y^{\star}$ and away from incorrect tokens. Consequently, under any fixed $\pi_{w}$ (with $w\geq 0$), $\mathbb{E}_{a\sim\pi_{w}}\Delta\ell[a]<\Delta\ell[y^{\star}]$. Taking expectation over data (and model stochasticity) yields the stated inequality. ∎

### v3

@@ Appendix C Discussion
Potential algorithmic improvements. A key limitation of the current implementation is the use of a fixed number of supervision steps for every ARC task. However, task complexity varies significantly; some tasks may benefit from a higher number of denoising steps, while others require fewer. This observation aligns with findings from the original TRM paper, which highlighted the significant contribution of the halting mechanism to the final performance. Therefore, explicitly predicting the necessary number of denoising steps for each task could potentially improve overall model efficiency and accuracy.
@@
Another promising direction for technical improvement involves the adoption of a discrete latent space. This approach has been successfully utilized in deep learning architectures such as the Dreamer model (Hafner et al., 2019) and VQ-VAE (Razavi et al., 2019), where latent spaces have proven to be robust and scalable for generative tasks.
@@
Alternative Improvement Generators. As described in §4.2, there are several viable methods to generate intermediate steps. Although the prior discrete diffusion serves as the main source in this work, our framework is designed to support various step-by-step approaches.